OpenStudy (anonymous):

\[(\sqrt{5x+5})^{2}=0\]

OpenStudy (jdoe0001):

\(\bf (\sqrt{5x+5})^{2}=0\\ -------------\\ \color{blue}{(\sqrt{5x+5})^{2}\implies \sqrt{(5x+5)^2}}\\ -------------\\ (\sqrt{5x+5})^{2}=0 \implies \sqrt{(5x+5)^2} = 0\)

OpenStudy (anonymous):

Am I supposed to take it out of radicals?

OpenStudy (anonymous):

Substitue a number under the radical instead of the expression that is there. 25 would be a good choice. Take a square root and then square the number and see what you get. Do the same with another perfect square and look for the pattern.

OpenStudy (anonymous):

Thank you again!

OpenStudy (anonymous):

You're Welcome again! Great job!